# STUPP-24-271

### v1

@@ Abstract
In this letter, we perform an almost general analysis of flavor mixing matrices to investigate the discriminative power of CP phases by the next-generation neutrino oscillation experiments. As an assumption, we neglect the 1-3 mixing of diagonalization for more hierarchical fermions $u,e$. Thus there are two sources of CP violation in the flavor mixing matrices, the intrinsic CP phase $\delta_{d,\nu}$ in diagonalization of less hierarchical fermions $d,\nu$ and relative phases between two unitary matrices.
By eliminating unphysical phases and imposing constraints of the three measured mixing angles, the flavor mixing matrices are analytically displayed by two phases and the 1-2 mixing $s_{u,e}$ of more hierarchical fermions. For sufficiently small 1-2 mixing $s_{e}$ of charged leptons, the Dirac phase $\delta$ is almost identical to the intrinsic phase of neutrinos $\delta_{\nu}$. Therefore, future detection of the Dirac phase by the Hyper-Kamiokande experiment indicates the observation of $\delta_{\nu}$. On the other hand, if such a CP violation is not observed, an upper limit is placed on a combination of $\delta_{\nu}$ and relative phases.
@@ 1 Introduction
In this current observational circumstance, a question arises; what kind of restrictions are generally placed on CP phases in the lepton sector, when $\delta$ is or is not observed by these experiments? For the CKM matrix, general analyses exist for hierarchical Yukawa matrices [5, 6] and many flavor texture studies [7]. However, no general treatment for each unitary matrix has been found for the two flavor mixing matrices $V_{\rm CKM}=U_{u}^{\dagger}U_{d}$ and $U_{\rm MNS}=U_{e}^{\dagger}U_{\nu}$. Therefore, in this letter, we analyze the quark and lepton mixing matrices in an almost general form for hierarchical mass matrices of the standard model fermions and discuss the distinguishability of CP phases of leptons by the next-generation experiments.
2 An almost general representation of flavor mixing matrices for hierarchical Yukawa structure
This section introduces an almost general form of flavor mixing matrices that comes from the hierarchical Yukawa matrices. At first, we define that a Yukawa matrix $Y_{f}$ of the Standard Model fermions $f=u,d,\nu,e$ is hierarchical if its elements satisfy $Y_{fii}\gg Y_{fij},Y_{ji}$ for $i>j$. Namely, for a matrix,
@@ 2 An almost general representation of flavor mixing matrices for hierarchical Yukawa structure
its elements $Y_{fij}$ satisfy
$Y_{f33}\gg Y_{f23},Y_{f32},Y_{f13},Y_{f31}\,,~{}Y_{f22}\gg Y_{f21},Y_{f12}\,.$ (2)
@@
Such $Y_{f}$ is easily realized by hierarchical Higgs vacuum expectation values (vevs) with appropriate chiral charges. Specifically, we consider a situation where the Yukawa matrices are generated by spontaneous symmetry breaking of flavon fields $\phi_{i}$ and $\tilde{\phi}_{j}$ that have chiral charges associated with the left-handed $i$ and right-handed $j$ generations. At the leading-order, Yukawa matrices are generated as
@@
Here, $y_{fij}$ denotes coupling constants between flavons and fermions $f$, and $\Lambda$ is some cutoff scale. Approximate chiral symmetries exist if these vevs satisfy $\langle{\phi_{3}}\rangle\gg\langle{\phi_{2}}\rangle\gg\langle{\phi_{1}}\rangle$ and $\langle{\tilde{\phi}_{3}}\rangle\gg\langle{\tilde{\phi}_{2}}\rangle\gg\langle{\tilde{\phi}_{1}}\rangle$. Various mass matrix models exhibit these properties.
2.1 An assumption from hierarchy of Yukawa matrices
For the aforementioned hierarchical Yukawa matrices, this letter deals with two flavor mixing matrices, the CKM matrix $V_{\rm CKM}$ and the MNS matrix $U_{\rm MNS}$.
@@ 2.1 An assumption from hierarchy of Yukawa matrices
Here, $U_{f}$ denote left-handed unitary matrices that diagonalize $Y_{f}$ and the neutrino mass matrix $m_{\nu}$. This $U_{f}$ is generally displayed in the PDG parameterization and its mixing angle $s_{ij}^{f}$.
To make the analysis concise, the following assumption is imposed.
Assumption:
We neglect the 1-3 mixing of $U_{f}$ for fermions $f$ with a stronger hierarchy of singular values (i.e. $f=u,e$),
This is justified by the approximate chiral symmetries described above, because the 1-3 mixing of $U_{u,e}$ is suppressed by $m_{f1}/m_{f2},m_{f1}/m_{f3}$.
@@
    $|\epsilon_{R}^{f}|\sim 1\gg|\epsilon_{L}^{f}|$: The 1-3 mixings are approximately evaluated as
@@
    We can neglect them because these mixings are sufficiently smaller than the 1-3 elements of the flavor mixing matrices $|(V_{\rm CKM})_{13}|\sim 10^{-3},|(U_{\rm MNS})_{13}|\sim 10^{2}$. Besides, for too large $\epsilon_{R}^{f}\gtrsim 0.1$, there is no longer an approximate symmetry.
@@
    Therefore, its contribution for leptons is at most a $\pm$10change in $s_{13}$. On the other hand, for quarks, the 1-3 mixing can have a large impact on $|V_{ub}|\simeq 0.003$. However, a similar hierarchy for down-type quarks leads to $\sqrt{m_{d1}/m_{d3}}\simeq 0.03$, and such a model with large 1-3 mixings is excluded unless a fine-tuned cancellation.
@@
    $|\epsilon_{L}|\gg|\epsilon_{R}|$: In this case, it is also hard to consider that there is an approximate $U(1)_{L}$ symmetry. The assumption does not hold and the unitary matrix for left-handed fields has large mixings. With the corresponding CKM matrix element $|V_{ub}|\simeq 0.003$ in mind, this situation seems difficult to realize in a unified theory.
@@
The assumption eliminates the intrinsic CP violation associated with $U_{u,e}$, because three generations are no longer involved in flavor mixing for the more hierarchical fermions. Thus, in this analysis, there are two sources of CP violation: relative phases between the two unitary matrices and intrinsic CP phase for the non-hierarchical fermions.
@@
In this letter, we will proceed with the analysis by considering $s_{12}^{e,u}\lesssim 0.1$.
2.2 Representation of flavor mixing matrices
Using this assumption, we organize the representation of the MNS (and CKM) matrix. For $U_{\nu,e}$ displayed in the PDG parameterization, the determinant of 2-3 submatrix can be made to unity by a redefinition of the overall phase. Then the MNS matrix is written by
$U_{\rm MNS}$ $\equiv O_{e}^{T}\begin{pmatrix}e^{i\rho}&0&0\\ 0&e^{i\sigma}&0\\ 0&0&e^{-i\sigma}\\ \end{pmatrix}O_{\nu}P,~{}~{}~{}O_{e}^{T}\equiv\begin{pmatrix}c_{e}&-s_{e}&0\\ s_{e}&c_{e}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}1&0&0\\ 0&c_{\tau}&-s_{\tau}\\ 0&s_{\tau}&c_{\tau}\\ \end{pmatrix}\,,$
$O_{\nu}P$ $\equiv\begin{pmatrix}1&0&0\\ 0&c_{\nu}&s_{\nu}\\ 0&-s_{\nu}&c_{\nu}\\ \end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{-i\delta_{\nu}}\\ 0&1&0\\ -s_{13}e^{i\delta_{\nu}}&0&c_{13}\\ \end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}e^{i\phi_{1}}&0&0\\ 0&e^{i\phi_{2}}&0\\ 0&0&e^{i\phi_{3}}\\ \end{pmatrix}.$ (13)
Here $s_{f}\equiv\sin f\,,\,c_{f}\equiv\cos f\,,\,s_{ij}\equiv\sin\theta_{ij}\,,\,c_{ij}\equiv\cos\theta_{ij}$ and $\rho,\sigma,\delta_{\nu},\phi_{i}$ are arbitrary phases. Since the phases $\phi_{i}$ only affect the Majorana phases, they are omitted hereafter.
@@ 2.2 Representation of flavor mixing matrices
Relations between these arguments $\chi,\eta$ and the phases are
@@
Therefore, the following two phases $\alpha$ and $\beta$ are physical.
@@
Finally, we discuss sign degrees of freedom for mixing angles. Continuous parameters in this notation are $s_{e},s_{ij}$ and two phases. Similar to the PDG parameterization, phase redefinitions of the six leptons can eliminate five signs. Since the sign of $s_{13}$ can be absorbed into the phase $e^{\pm i\beta}$, the six $s_{ij}$ and $c_{ij}$ are set to be positive. By multiplying diag $(-1,-1,1)$ from the left, the sign of $c_{e}$ is inverted without changing the sign of $s_{ij}$ and $c_{ij}$. Therefore, only the sign of $s_{e}$ is physical because the sign of the two $s_{e}$ cannot be absorbed into the phase $\alpha$. We can specifically check that the physical sign in the CP violation is sign$(c_{e}s_{e})$ later. Next subsection shows that there are three free parameters $\alpha,\beta,s_{e}$ in the MNS matrix, and two in the CKM matrix because the CP phase is further determined.
@@ 2.3 Constraints on mixing angles
We will consider how the observables constrain these parameters. The best-fit values of the latest global fit [9] are used as input parameters for the MNS matrix. The three mixing angles of the normal hierarchy (NH) without Super-Kamiokande (SK) are
@@
The reason for using NH without SK is that the values of inverted hierarchy (IH) with or without SK are close to these values. Although the inclusion of the SK data makes $s_{23}^{\rm PDG}$ about $0.1$ smaller for NH, the qualitative behavior in the following discussion remains the same.
Three mixing parameters $s_{ij}$ are constrained from these input parameters. Hereafter, the above three input parameters displayed in PDG parameterization are denoted as $\sin\theta_{ij}^{\rm PDG}\equiv S_{ij}\,,\cos\theta_{ij}^{\rm PDG}\equiv C_{ij}$. First, the 3-3 element of $U_{\rm MNS}$, which is the same as the PDG parameterization, determines $s_{23}$;
@@
$s_{13}^{2}$ $={A\pm\sqrt{B}\over(c_{e}^{4}+2c_{e}^{2}s_{e}^{2}\cos 2(\alpha-\beta)+s_{e}^{4})}\,,$ (25)
@@
For sufficiently small $s_{13}$, since the two independent solutions are absorbed to the sign of $\cos(\alpha-\beta)$, the sign $A+\sqrt{B}$ is chosen as the solution that reproduces the correct $U_{13}$. In this case, the other solution corresponds to the dashed line in Fig. 1.
@@
This result is easy to understand because the mixing matrix of $s_{ij}$ is a “reverse rotation” of the PDG representation of $S_{ij}$. From this, there exists a critical value $s_{e}^{c}$ of $s_{e}$ such that $s_{13}=0$.
@@
This value is about 0.2 for the MNS matrix (and 0.09 for the CKM matrix). For larger $s_{e}>s_{e}^{c}$, as can be seen in Fig. 1, the two solutions in Eq. (25) (with different signs before the root sign) are physically inequivalent.
@@
By substituting the solutions of $s_{23}$ (21) and $s_{13}$ (25) to the constraint $|U_{12}|^{2}=C_{13}^{2}S_{12}^{2}$, $s_{12}^{2}$ is similarly obtained by solving a quadratic equation. As in the case of $s_{13}$, the sign before the root sign can be absorbed into the term $2c_{23}s_{e}(c_{13}c_{e}\cos\alpha+s_{13}s_{23}s_{e}\cos\beta)$.
@@
Since this constraint has a similar structure to Eq. (23), we choose $s_{12}\simeq A^{\prime}+\cos\alpha\,B^{\prime}>0$ as a positive solution that reproduces the matrix elements of $U_{\rm MNS}$. Here, $A^{\prime},B^{\prime}$ are functions of several trigonometric functions. We will not discuss the critical value of $s_{12}$ because $s_{12}$ is relatively large for both the CKM and MNS matrices.
By imposing the three constraints, we can treat CP violation as a function of $\alpha,\beta$ and $s_{e}$. The rough parameter dependences of the three solutions are as follows.
  • $s_{23}$: From Eq. (21), dependences of $\alpha,\beta$ on $s_{23}$ are relatively small because these phases enter in the second order of $s_{e}$ and $s_{13}$.
@@
  • $s_{13}$: From Eq. (23), $s_{13}$ depends on $\alpha-\beta$. The expansion for small $s_{e}$ is
@@
  • $s_{12}$: Since $\beta$ dependence is suppressed by $s_{13}s_{e}$ in Eq. (30), $s_{12}$ has an almost only $\alpha$ dependence. Therefore, the behavior of the solution is

### v2

@@ Abstract
In this letter, we perform an almost general analysis of flavor-mixing matrices $V_{\rm CKM}$ and $U_{\rm MNS}$ to investigate the discriminative power of CP phases by next-generation neutrino oscillation experiments. As an approximation, we neglect the 1-3 mixing of diagonalization for more hierarchical fermions $u,e$. Thus there are two sources of CP violation in $V_{\rm CKM}$ and $U_{\rm MNS}$, the intrinsic CP phase $\delta_{d,\nu}$ in diagonalization of less hierarchical fermions $d,\nu$ and relative phases between two unitary matrices. By eliminating unphysical phases and imposing constraints of the three measured mixing angles, the flavor-mixing matrices are analytically displayed by two phases and the 1-2 mixing $s_{u,e}$ of more hierarchical fermions. For sufficiently small 1-2 mixing $s_{e}$ of charged leptons, the Dirac phase $\delta$ is mostly identical to the intrinsic phase of neutrinos $\delta_{\nu}$. Therefore, future detection of the Dirac phase indicates the observation of $\delta_{\nu}$. On the other hand, if such a CP violation is not observed, an upper limit is placed on a combination of $\delta_{\nu}$ and relative phases.
@@ 1 Introduction
In this current observational circumstance, a question arises; what kind of restrictions are generally placed on CP phases in the lepton sector, when $\delta$ is or is not observed by these experiments? For the CKM matrix, general analyses exist for hierarchical Yukawa matrices [5, 6] and many flavor texture studies [7]. Phenomenological analyses have also been studied for the next-generation neutrino oscillation experiments [8, 9, 10, 11] and the Dirac phase [12, 13, 14, 15, 16, 17]. However, no general treatment for each unitary matrix has been found for the two flavor-mixing matrices $V_{\rm CKM}=U_{u}^{\dagger}U_{d}$ and $U_{\rm MNS}=U_{e}^{\dagger}U_{\nu}$. Therefore, in this letter, we analyze the mixing matrices of quarks and leptons in an almost general form for hierarchical mass structures of the fermions and discuss the distinguishability of CP phases of leptons by the next-generation experiments.
2 An almost general representation of flavor-mixing matrices for hierarchical Yukawa structure
This section introduces an almost general form of flavor-mixing matrices that originate from hierarchical Yukawa matrices. At first, we define that a Yukawa matrix $Y_{f}$ of the Standard Model fermions $f=u,d,\nu,e$ is hierarchical if its elements satisfy $|Y_{fii}|\gg|Y_{fij}|,|Y_{ji}|$ for $i>j$. Namely, for a matrix,
@@ 2 An almost general representation of flavor-mixing matrices for hierarchical Yukawa structure
absolute values of its elements $|Y_{fij}|$ satisfy
$|Y_{f33}|\gg|Y_{f23}|,|Y_{f32}|,|Y_{f13}|,|Y_{f31}|\,,~{}{\rm and}~{}|Y_{f22}|\gg|Y_{f21}|,|Y_{f12}|\,.$ (2)
@@
Such $Y_{f}$ is easily realized by hierarchical vacuum expectation values (vevs) of scalar (flavon) fields with appropriate chiral charges. Specifically, we consider a situation where the Yukawa matrices are generated by vevs of flavon fields $\phi_{i}$ and $\tilde{\phi}_{j}$ that have chiral charges associated with the left-handed $i$ and right-handed $j$ generations. At leading order, Yukawa matrices are generated as
@@
Here, $y_{fij}$ denotes coupling constants between flavons and fermions $f$, and $\Lambda$ is some cutoff scale. Approximate chiral symmetries exist if these vevs satisfy $|\langle{\phi_{3}}\rangle|\gg|\langle{\phi_{2}}\rangle|\gg|\langle{\phi_{1}}\rangle|$ and $|\langle{\tilde{\phi}_{3}}\rangle|\gg|\langle{\tilde{\phi}_{2}}\rangle|\gg|\langle{\tilde{\phi}_{1}}\rangle|$. Various models of mass matrices exhibit these properties.
2.1 An approximation from hierarchy of Yukawa matrices
For the hierarchical Yukawa matrices, this letter deals with two flavor-mixing matrices, the CKM matrix $V_{\rm CKM}$ and the MNS matrix $U_{\rm MNS}$.
@@ 2.1 An approximation from hierarchy of Yukawa matrices
Here, $U_{f}$ denote left-handed unitary matrices that diagonalize $Y_{f}$ and the mass matrix of neutrinos $m_{\nu}$. This $U_{f}$ is generally displayed by the PDG parameterization as
$U_{f}=\Phi_{L}^{f}U_{f}^{0}\Phi_{R}^{f}\,,$ (6)
with diagonal phase matrices $\Phi^{f}_{L,R}$. The matrix $U_{f}^{0}$ is a function $U_{f}^{0}=f(s_{12}^{f},s_{23}^{f},s_{13}^{f},\delta^{f})$ of mixing angles $s_{ij}^{f}$ and the intrinsic CP phase $\delta^{f}$.
To make the analysis concise, the following approximation is imposed.
Approximation:
We neglect the 1-3 mixing of $U_{f}$ for fermions $f$ with a stronger hierarchy of singular values (i.e. $f=u,e$).
Justification:
The 1-3 mixing of $U_{f}$ is suppressed by $m_{f1}/m_{f3}$ (or a power of the ratio) by the approximate chiral symmetry described above.
@@
    $|\epsilon_{R}^{f}|\sim 1\gg|\epsilon_{L}^{f}|$: From $(Y_{f})_{11}/(Y_{f})_{33}\simeq m_{f1}/m_{f3}$, the 1-3 mixings are approximately evaluated as
@@
    These mixings are safely neglected because they are sufficiently smaller than the 1-3 elements of the flavor-mixing matrices $|(V_{\rm CKM})_{13}|\sim 10^{-3},|(U_{\rm MNS})_{13}|\sim 10^{-2}$. Besides, for too large $\epsilon_{R}^{f}\gtrsim 0.1$, there is no longer the approximate symmetry for right-handed fields.
@@
    Its contribution for leptons is at most a $\pm$10change in $U_{e3}$. On the other hand, the 1-3 mixing for quarks can have a large impact on $|V_{ub}|\simeq 0.003$. However, a similar hierarchy for down-type quarks leads to $\sqrt{m_{d1}/m_{d3}}\simeq 0.03$, and such a model with large 1-3 mixings is excluded unless a fine-tuned cancellation.
@@
    $|\epsilon_{L}|\gg|\epsilon_{R}|$: In this case, it is also hard to consider that there is an approximate $U(1)_{L}$ symmetry. The approximation does not hold and the unitary matrix for left-handed fields has large mixings. With the corresponding CKM matrix element $|V_{ub}|\simeq 0.003$ in mind, this situation seems difficult to realize in a unified theory.
@@
In subsequent analyses, we consider parameter regions of $s_{12}^{e,u}\lesssim 0.1$.
2.2 Representation of flavor-mixing matrices
Using this approximation, we organize an almost general representation of the flavor-mixing matrices. For unitary matrices $U_{\nu,e}=\Phi_{L}^{\nu,e}U_{\nu,e}^{0}\Phi_{R}^{\nu,e}$ displayed in the PDG parameterization, there is relative phases $\Phi_{L}^{e\dagger}\Phi_{L}^{\nu}=$ diag $(e^{i\rho},e^{i\sigma},e^{i\sigma^{\prime}})$ between $U_{\nu}^{0}$ and $U_{e}^{0}$. By a redefinition of the overall phase, we define $\sigma^{\prime}\equiv-\sigma$ to set the determinant of the 2-3 submatrix to unity. Then the MNS matrix is written by
$U_{\rm MNS}$ $\equiv\Phi_{R}^{e\dagger}U_{e}^{0\dagger}\begin{pmatrix}e^{i\rho}&0&0\\ 0&e^{i\sigma}&0\\ 0&0&e^{-i\sigma}\\ \end{pmatrix}U_{\nu}^{0}\Phi_{R}^{\nu},~{}~{}~{}U_{e}^{0\dagger}\equiv\begin{pmatrix}c_{e}&-s_{e}&0\\ s_{e}&c_{e}&0\\ 0&0&1\\ \end{pmatrix}\begin{pmatrix}1&0&0\\ 0&c_{\tau}&-s_{\tau}\\ 0&s_{\tau}&c_{\tau}\\ \end{pmatrix}\,,$
$U_{\nu}^{0}$ $\equiv\begin{pmatrix}1&0&0\\ 0&c_{\nu}&s_{\nu}\\ 0&-s_{\nu}&c_{\nu}\\ \end{pmatrix}\begin{pmatrix}c_{13}&0&s_{13}e^{-i\delta_{\nu}}\\ 0&1&0\\ -s_{13}e^{i\delta_{\nu}}&0&c_{13}\\ \end{pmatrix}\begin{pmatrix}c_{12}&s_{12}&0\\ -s_{12}&c_{12}&0\\ 0&0&1\\ \end{pmatrix}.$ (14)
Here $s_{f}\equiv\sin f\,,\,c_{f}\equiv\cos f\,,\,s_{ij}\equiv\sin\theta_{ij}\,,\,c_{ij}\equiv\cos\theta_{ij}$. Since the phases $\Phi_{R}^{\nu,e}$ do not affect the Dirac phase, they are omitted hereafter. The CKM matrix is also defined in the same way.
The approximation eliminates the intrinsic CP phase $\delta_{u,e}$ associated with $U_{u,e}$, because three generations are no longer involved in flavor-mixing for the more hierarchical fermions. Thus, in this analysis, there are two sources of CP violation: relative phases between the two unitary matrices $\rho,\sigma$ and intrinsic CP phase $\delta_{d,\nu}$ for the less hierarchical fermions.
@@ 2.2 Representation of flavor-mixing matrices
Relations between these arguments $\chi,\eta$ and parameters $\sigma,\nu,\tau$ are
@@
Therefore, the following two phases $\alpha$ and $\beta$ have physical significance.
@@
Finally, we discuss sign degrees of freedom for mixing angles. Continuous parameters in this notation are $s_{e},s_{ij}$ and two phases $\alpha,\beta$. Similar to the PDG parameterization, phase redefinitions of six fermions can eliminate five signs. Since the sign of $s_{13}$ can be absorbed into the phase $e^{\pm i\beta}$, the six $s_{ij}$ and $c_{ij}$ are set to be positive. By multiplying diag $(-1,-1,1)$ from the left of Eq. (19), the sign of $c_{e}$ is inverted without changing the sign of $s_{ij}$ and $c_{ij}$. As a result, only the sign of $s_{e}$ is physical because the sign of two $s_{e}$ cannot be absorbed into the phase $\alpha$. Later, we specifically check that the physical sign of the CP violation is sign$(c_{e}s_{e})$.
@@ 2.3 Constraints on mixing angles
The six continuous parameters ($s_{e},s_{ij},\alpha$ and $\beta$) are constrained by the three observed mixing angles, resulting in three free parameters $\alpha,\beta,s_{e}$ in the MNS matrix. For this purpose, the best-fit values of the latest global fit [19] are used as input parameters for the MNS matrix. The three mixing angles of the normal hierarchy (NH) without Super-Kamiokande (SK) are
@@
The reason for using NH without SK is that the values of inverted hierarchy (IH) with or without SK are close to these values. Although the inclusion of the SK data makes $\theta_{23}^{\rm NH}$ about $0.1$ smaller, the qualitative behavior remains the same in the following discussion.
Three mixing parameters $s_{ij}$ are constrained from $\theta_{ij}^{\rm NH}$. Hereafter, the above three input parameters are denoted as $\sin\theta_{ij}^{\rm NH}\equiv S_{ij}\,,\cos\theta_{ij}^{\rm NH}\equiv C_{ij}$. First, the 3-3 element of $U_{\rm MNS}$, which is the same as the PDG parameterization, determines $s_{23}$;
@@
$s_{13}^{2}$ $={A\pm B\over(c_{e}^{4}+2c_{e}^{2}s_{e}^{2}\cos 2(\alpha-\beta)+s_{e}^{4})}\,,$ (26)
@@
For sufficiently small $s_{13}$, the solution $A+B$ is chosen because it reproduces the correct $U_{13}$. The other solution corresponds to the dashed line in Fig. 1, because the sign is identified with the sign of $\cos(\alpha-\beta)$.
@@
This result is easy to understand because the mixing matrix of $s_{ij}$ is a “reverse rotation by $s_{e}$” of the PDG representation of $S_{ij}$. From this, there exists a critical value $s_{e}^{c}$ of $s_{e}$ such that $s_{13}=0$.
@@
This value is about 0.2 for the MNS matrix (and 0.09 for the CKM matrix). For larger $s_{e}>s_{e}^{c}$, as can be seen in Fig. 1, the two solutions in Eq. (26) are physically inequivalent.
@@
By substituting the solutions of $s_{23}$ (22) and $s_{13}$ (26) to the constraint $|U_{12}|^{2}=C_{13}^{2}S_{12}^{2}$, $s_{12}^{2}$ is similarly obtained by solving a quadratic equation. As in the case of $s_{13}$, the sign before the root sign can be absorbed into a term $2c_{23}s_{e}(c_{13}c_{e}\cos\alpha+s_{13}s_{23}s_{e}\cos\beta)$.
@@
Since this constraint has a similar structure to Eq. (24), we choose $s_{12}^{2}\simeq A^{\prime}+\cos\alpha\,B^{\prime}>0$ as a positive solution that reproduces the matrix elements of $U_{\rm MNS}$. Here, $A^{\prime},B^{\prime}$ are functions of several trigonometric functions. We will not discuss the critical value of $s_{12}$ because it is relatively large for both the CKM and MNS matrices.
Imposing the three constraints, we can treat the CP violation in flavor-mixing matrices as a function of $\alpha,\beta$ and $s_{e}$. Rough parameter dependences of the three solutions are as follows.
  • $s_{23}$: From Eq. (22), dependences of $\alpha,\beta$ on $s_{23}$ are relatively small because these phases depend on the second order of $s_{e}$ and $s_{13}$.
@@
  • $s_{13}$: From Eq. (23), $s_{13}$ depends on $\alpha-\beta$. An expansion of $s_{13}$ for small $s_{e}$ is
@@
  • $s_{12}$: Since dependence of $\beta$ is suppressed by $s_{13}s_{e}$ in Eq. (31), $s_{12}$ has an almost only $\alpha$ dependence for a small $s_{e}$. Therefore, the behavior of the solution is